# Constructing Per-Shot Bitrate Ladders using Visual Information Fidelity

### v1

@@ VI Conclusion and Future Work
We studied the efficacy of using features drawn from Visual Information Fidelity, extracted over scales and subbands, as features for constructing video bitrate and quality ladders. We proposed multiple feature sets based on VIF features, as well as an ensemble of VIF features and low-level features for constructing content-gnostic ladders for video delivery. The advantage of the VIF feature sets is that they can be easily calculated using a full-reference (FR) quality estimation model (e.g., VMAF) that is already widely deployed in video-delivery pipelines. We designed multiple regression models trained on these video content-gnostic features, along with metadata such as frame width and height, and bitrate or quality, to predict the quality or bitrate of compressed videos, respectively. We deployed these quality and bitrate prediction models to construct per-shot bitrate ladders with bitrate steps or quality ladders with VMAF steps. We compared our proposed models against existing popular techniques, including predicting cross-over bitrates and quality using low-level features. We also designed counterpart models that predict cross-over VMAF points and bitrates. Our observations indicated that ladders constructed by predicting cross-over points generally gave the best performance. When considering various quality and bitrate prediction methodologies, the use of an ensemble of low-level features and VIF features to predict quality ($\text{LLF}_{2}$, $\text{VIFF}_{7}$) in the construction of bitrate ladders, and the prediction of bitrates ($\text{LLF}_{3}$, $\text{VIFF}_{9}$) in the construction of quality ladders, obtained excellent performance against fixed bitrate ladder and reference bitrate ladders constructed from exhaustive encoding. The performances of these methods is close to, if not superior in some cases, to methods that predicted cross-over points. These results suggest a significant potential for constructing more efficient per-shot bitrate and quality ladders. In the future, we aim to explore the performances of these methods across various presets and codecs.

### v2

@@ V Conclusion and Future Work
We studied multiple feature sets based on VIF features and an ensemble of VIF features and low-level features, for constructing content-gnostic ladders. Our proposed methods predict per-shot bitrate or quality ladders of a video using regressors trained to predict video quality or bitrate of compressed videos using video features and metadata, respectively. We compared our proposed models against existing popular techniques, including predicting cross-over bitrates and quality using low-level features. Our observations show that regressors trained on low-level features and VIF features demonstrated significant gains in bitrate and quality against the fixed bitrate ladder and small losses against the reference bitrate ladder. Their performance is followed closely by the cross-over point prediction models, regressors trained on low-level features and VIF features, respectively. Our methods predict approximate bitrates and video quality without compression or quality estimation, allowing us to construct bitrate and quality ladders, without relying on exhaustive encoding. These results suggest a significant potential for efficiently constructing per-shot bitrate and quality ladders.
Future Work
In the future, we aim to conduct experiments on larger datasets to better understand and improve per-shot ladder construction techniques. We intend to evaluate the efficacy of these methods across diverse presets and codecs. Furthermore, we will investigate the transferability of constructed per-shot bitrate ladders across multiple encoder settings.